# Breaking the Low-Rank Dilemma of Linear Attention

### v3

@@ 4.1 Image Classification
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification. We use ”CNN” to refer to convolutional neural networks, ”Trans” to refer to Vision Transformers, and ”Linear” to refer to models based on linear operators.
@@ Results.
4.2 Inference Speed
4.3 Object Detection and Instance Segmentation
Settings.
We adopt MMDetection [5] to implement RetinaNet [42], Mask-RCNN [29] and Cascade Mask R-CNN [3]. Following previous works [19, 45, 58], we use the commonly used “1×1$\times$1 ×” (12 training epochs) setting for the RetinaNet and Mask R-CNN. Besides, we use “$3\times+\mathrm{MS}$” for Mask R-CNN and Cascade Mask R-CNN.
@@ 4.3 Object Detection and Instance Segmentation
Table 4: Comparison to other backbones using Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
@@
Table 5: Comparison to other backbones using Cascade Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
@@
Table 6: Comparison to other backbones using RetinaNet and Mask R-CNN with “1×1$\times$1 ×” schedule.
@@ 4.6 Ablation Study
In this section, we conduct ablation experiments on various key modules of the model to verify the performance enhancements contributed by each component.
@@ Appendix A More Analysis About $\alpha$
Figure 9: Overall architecture of RAVLT.
Appendix C Implementation Details
Overall Architectures.
Similar to previous works [19, 35, 18], RAVLT employs a four-stage architectural design, as shown in Fig. 9.
@@ Appendix C Implementation Details
We apply RetinaNet [42], Mask-RCNN [29], and Cascaded Mask R-CNN [3] as the frameworks based on the MMDetection [5] to evaluate our models. The models are trained under “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings. For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS”, multi-scale training strategy is applied to randomly resize the shorter side between 480 to 800 pixels. We use the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. For Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.

### v5

@@ 3.2 Rank-Augmented Linear Attention
It should be emphasized that the summation of matrices can indeed increase their rank. This is fundamentally different from a simple linear combination, which cannot increase the rank. For example:
$1\times\begin{bmatrix}1&1&1\\ 2&2&2\\ 3&3&3\end{bmatrix}+1\times\begin{bmatrix}1&2&3\\ 1&2&3\\ 1&2&3\end{bmatrix}=\begin{bmatrix}2&3&4\\ 3&4&5\\ 4&5&6\end{bmatrix}$
A linear combination of two rank-1 matrices can result in a matrix with rank 3.
@@ 4.1 Image Classification
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification. We use "CNN" to refer to convolutional neural networks, "Trans" to refer to Vision Transformers, and "Linear" to refer to models based on linear operators.
@@ Results.
Table 4: Comparison to other backbones using Mask R-CNN with "$3\times+\mathrm{MS}$" schedule.
4.2 Inference Speed
4.3 Object Detection and Instance Segmentation
Settings.
We adopt MMDetection [5] to implement RetinaNet [41], Mask-RCNN [28] and Cascade Mask R-CNN [3]. Following previous works [18, 44, 57], we use the commonly used “1×1$\times$1 ×" (12 training epochs) setting for the RetinaNet and Mask R-CNN. Besides, we use “$3\times+\mathrm{MS}$" for Mask R-CNN and Cascade Mask R-CNN.
@@ 4.3 Object Detection and Instance Segmentation
Table 5: Comparison to other backbones using Cascade Mask R-CNN with "$3\times+\mathrm{MS}$" schedule.
@@
Table 6: Comparison to other backbones using RetinaNet and Mask R-CNN with “1×1$\times$1 ×" schedule.
@@ 5 Conclusion
6 Acknowledgements
This work is partially funded by Beijing Natural Science Foundation (4252054), Youth Innovation Promotion Association CAS(Grant No.2022132), Beijing Nova Program(20230484276), and CCF-Kuaishou Large Model Explorer Fund (NO. CCF-KuaiShou 2024005).
[…]
Supplementary Material
@@ Appendix A More Analysis About $\alpha$
Appendix C Experimental Details
@@
We apply RetinaNet [41], Mask-RCNN [28], and Cascaded Mask R-CNN [3] as the frameworks based on the MMDetection [5] to evaluate our models. The models are trained under “1 $\times$" (12 training epochs) and “3 $\times$ +MS" (36 training epochs with multi-scale training) settings. For the “1 $\times$" setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS", multi-scale training strategy is applied to randomly resize the shorter side between 480 to 800 pixels. We use the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. For Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.